OpenStudy (anonymous):

The midpoint of UV is (5,-11). One of the coordinates are U(3,5). Find V

OpenStudy (freckles):

\[\text{ the midpoint of the line segment from point } (a,b) \text{ \to } (c,d) \text{ is } \\ (\frac{a+c}{2},\frac{b+d}{2})\]

OpenStudy (freckles):

you are given: \[\frac{3+c}{2} =5 \text{ and } \frac{5+d}{2}=-11\]

OpenStudy (freckles):

where (c,d) is point V

OpenStudy (anonymous):

Would I have to plug something into the formula? How would the end result look

OpenStudy (freckles):

do you know how to solve equations yet?

OpenStudy (freckles):

maybe not... I guess since you are asking how to solve (3+c)/2=5 and also solve (5+d)/2=-11

OpenStudy (freckles):

the first I would do would be to multiply both sides by 2

OpenStudy (freckles):

for both equations

OpenStudy (anonymous):

Ok so that would give 5+d=-22?

OpenStudy (freckles):

yes for the second equation you have d+5=-22

OpenStudy (anonymous):

and then 3+c=10

OpenStudy (freckles):

do you the +5 on the side with d on it you are going to minus 5 on both sides because you know 5-5=0 so you will have d by itself

OpenStudy (freckles):

and that is right for the first equation so far

OpenStudy (anonymous):

And then you do the same to the other one too?

OpenStudy (freckles):

well not subtract 5 on both sides but subtract 3 on both sides yes

OpenStudy (anonymous):

So d=-27 and c=7?

OpenStudy (freckles):

ok great now remember (c,d) is the point V so your answer would be (7,-27) is the other endpoint V

OpenStudy (freckles):

so the midpoint of (7,-27) and (3,5) is \[(\frac{3+7}{2},\frac{-27+5}{2})=(\frac{10}{2},\frac{-22}{2})=(5,-11)\]

OpenStudy (anonymous):

Ok great I get it now

OpenStudy (freckles):

cool stuff have a good day :)

OpenStudy (anonymous):

Thank you very much for your help! My sister doesn't know how to do this, thank you

OpenStudy (freckles):

well go teach her then

OpenStudy (anonymous):

Ok i will now